# GlobalMood: A cross-cultural benchmark for music emotion recognition

### v1

@@ Abstract
Human annotations of mood in music are essential for music generation and recommender systems. However, existing datasets predominantly focus on Western songs with mood terms derived from English, which may limit generalizability across diverse linguistic and cultural backgrounds. To address this, we introduce ‘GlobalMood’, a novel cross-cultural benchmark dataset comprising 1,180 songs sampled from 59 countries, with large-scale annotations collected from 2,519 individuals across five culturally and linguistically distinct locations: U.S., France, Mexico, S. Korea, and Egypt. Rather than imposing predefined mood categories, we implement a bottom-up, participant-driven approach to organically elicit culturally specific music-related mood terms. We then recruit another pool of human participants to collect 988,925 ratings for these culture-specific descriptors. Our analysis confirms the presence of a valence-arousal structure shared across cultures, yet also reveals significant divergences in how certain mood terms, despite being dictionary equivalents, are perceived cross-culturally. State-of-the-art multimodal models benefit substantially from fine-tuning on our cross-culturally balanced dataset, as evidenced by improved alignment with human evaluations—particularly in non-English contexts. More broadly, our findings inform the ongoing debate on the universality versus cultural specificity of emotional descriptors, and our methodology can contribute to other multimodal and cross-lingual research.
@@ 1 Introduction
Figure 1: Elicitation and refinement of music mood terms through iterative participant chains. (A) Schematic illustration of the collaborative tagging process within a participant chain. Participants contribute new emotion-related word tags for each song, rate the relevance of existing tags, and can also flag irrelevant content, creating a dynamic refinement system. (B) The twenty most reliable emotion tags in each language, ranked by their tag scores. Y-axis labels display tags in their original language (left) and English translations (right).
@@
Music evokes diverse emotional responses in listeners, spanning a wide spectrum beyond basic emotional categories [1, 2]. A central challenge in Music Information Retrieval (MIR) is designing algorithms that can replicate this emotional sensitivity. This is crucial for building recommendation systems that align with listeners’ mood and context [3, 4], and for generating music that resonates with individual preferences [5]. More broadly, understanding how music conveys emotion is a core question in the science of music [6, 7]. To date, however, most algorithms have been trained on datasets derived from Western listeners and Western music, using mood taxonomies primarily based on English language (e.g., MIREX [8]).
@@
A significant challenge is creating cross-cultural models capable of handling non-Western music and mood vocabularies beyond English. Addressing this challenge is essential to developing algorithms that accurately reflect global users’ preferences, including those whose musical tastes extend beyond the limited range of styles currently represented in training datasets. Moreover, without capturing culturally specific mood terms, especially those difficult to translate, key aspects of musical meaning may be missed entirely. Direct dictionary translations of English terms may be insufficient, as mood terms are deeply cultural and may lack exact equivalents [9, 10, 11, 12].
@@
To address these issues, we introduce ‘GlobalMood’, a new benchmark dataset designed to support culturally inclusive and linguistically diverse mood recognition in music. Our contribution innovates along three key dimensions: (i) the diversity of musical stimuli, drawn from 59 countries; (ii) the diversity of annotators, spanning five distinct cultural and linguistic regions; (iii) a data-driven approach for collecting mood descriptors, generated organically by participants in their own language during the annotation process.
@@
Data were collected through two stages involving a total of 2,519 participants and 1,180 songs balanced evenly across 59 countries: In the first stage (Section 4.1; Figure 1), using a smaller subset of 200 songs, we employed a recently developed iterative task that combines open-ended elicitation with collective refinement [13, 14, 11]. Rather than asking listeners to choose from a fixed list of pre-defined mood terms, we asked them to describe the perceived mood conveyed by a piece of music using free-text tags in their native language, at the same time, rate the tags provided by previous listeners. This approach was key to uncovering mood terms that would otherwise be overlooked by predefined, English-based taxonomies (such as ’appeal/plead’ that appears in Korean only).
@@
In the second stage (Section 4.2; Figure 2), we selected the top 20 elicited terms per language and crowdsourced ratings for each tag across the entire set of 1,180 songs. This resulted in a total of 988,925 mood ratings, creating the most comprehensive open-source cross-cultural emotion annotation dataset in Music Emotion Recognition (MER) to date (available at: https://osf.io/8c5f2).
@@
We leveraged GlobalMood to test several recent multimodal and multilingual models (Gemini, CLAP) by evaluating their performance on 15-second audio clips under zero-shot, few-shot, and fine-tuned scenarios (Section 4.3; Figure 3). Our results demonstrate that models trained exclusively on English data underperform when applied to certain cultural contexts, but fine-tuning with our cross-cultural dataset significantly improves performance in non-English countries. This highlights the critical importance of cross-cultural data in both training MER models and establishing appropriate benchmarks for their evaluation.
@@
Figure 2: Association between mood terms across languages. (A) MDS visualization of mood terms based on mean ratings across the full song set. Terms positioned closer together exhibit similar rating patterns across songs, suggesting similar interpretations across languages. (B) Comparison of terms with direct translation equivalents across languages. The area size indicates the degree of semantic divergence in perception despite apparent translation equivalence.
@@ 2 Related Works
2.1 Music Mood Annotation Datasets
@@
Several datasets have been developed for MER systems with varying annotation approaches. Early examples include the widely used MIREX 2007 mood dataset [8] with 240-250 Western songs in five mood clusters derived from AllMusic’s English tags (e.g., ‘passionate–rousing’, ‘wistful–bittersweet’), and CAL500 [15] with 500 Western pop/rock songs annotated using 18 English mood terms by U.S. undergraduate listeners. Over time, larger datasets appeared: the DEAM corpus (MediaEval ‘Emotion in Music’ dataset [16]) containing 2,058 song excerpts with continuous valence/arousal annotations; mood tags mined from large corpora of Spotify music playlists [17]; and the MTG-Jamendo dataset [18], which provides mood/theme tags for 18,486 songs. Notably, Jamendo’s tags were freely crowdsourced (56 unique mood labels), which introduced more label variety but still almost entirely in English.
@@
A common limitation across these datasets is their reliance on predefined English descriptors, many of which stem from Western music psychology. For instance, the Geneva Emotional Music Scale (GEMS) defines 45 emotion descriptors (e.g., ‘joyful activation’) based on studies with European listeners [1], and this taxonomy has been used to annotate datasets like Emotify [19]. Similarly, the mood categories in MIREX and CAL500 were fixed in advance (drawn from AllMusic or prior literature) and presented to annotators as a closed set of options. Consequently, these top-down approaches restrict annotators to the moods the researchers envisioned, leaving any unlisted mood nuances uncaptured and undocumented.
@@
Acknowledging these limitations, recent research has begun exploring MER beyond the Western-centric scope. Hu et al. [20] examined mood annotations of K-pop songs provided by both Korean and American listeners. Their approach involved translating the original MIREX mood categories into Korean for local annotators (included in the 2014 MIREX challenge). Although this allowed direct comparisons of mood classification between Korean and American listeners, it inherently restricted Korean annotations to terms originally defiend within Western contexts.
@@
More recently, Lee et al. [10] compiled a balanced set of American, Brazilian, and Korean songs and gathered mood annotations across nine categories, where annotators rated songs both from their own and the other two countries. They showed that certain mood terms like ‘energetic’ and ‘sad’ are highly consistent across cultures, while more abstract concepts like ‘love’ and ‘dreamy’ diverge considerably. Similar findings have been reported by other studies [12, 11], highlighting that when mood descriptors are imposed from one language onto another, important meanings can simply be ‘lost in translation’.
@@
In summary, while existing MER datasets and research have laid a solid groundwork, they remain limited by insufficient linguistic and cultural diversity. Because many are predominantly English-based and rely on top-down annotation strategies, they may overlook how people in other cultural contexts perceive mood of the music. As awareness of these limitations grows, there have been increasing calls for more inclusive approaches. Our work responds to this need by introducing a crowd-sourcing approach that organically elicits mood terms across languages, provided with a diverse, globally balanced set of musical examples.
@@ 2.2 Audio LLMs: the New Frontier in Music Tagging
Recent advances in multimodal large language models (LLMs) have opened promising avenues for downstream MIR tasks, including mood recognition. These models combine the reasoning capabilities of LLMs with audio perception systems (audio LLMs), enabling more flexible and nuanced music understanding than traditional classification approaches [21, 22, 23].
Models like MuLan [24] and MERT [21] have demonstrated potential for zero-shot music mood classification by embedding audio and natural language descriptions in a shared semantic space. However, comprehensive benchmarks such as the MuChoMusic [22] highlight a crucial limitation: these models rely heavily on language modality and do not attend sufficiently to audio, often failing with more nuanced audio examples for downstream MIR tasks. This limitation could be particularly critical for non-Western music and non-English mood descriptors, given that their training data are largely from Western contexts.
@@ 3.1 Participants
We recruited two independent sets of participants across the two stages of our data collection: Stage 1 for emotion term elicitation (N = 778; see Section 4.1) and Stage 2 for subsequent ratings on top 20 terms (N = 1,741; see Section 4.2). Participants had to be at least 18 years old, reside in the target country, and speak the target language as their primary language. Participants from the US were recruited through Prolific, while participants from the other four countries (France, Mexico, S. Korea, and Egypt) were recruited through the CINT platform. All participants provided informed consent under an approved protocol (see Section 6). Participants were instructed to wear headphones and take part in the study in a quiet environment. They had to pass a headphone screening task [26] and a language proficiency test [27] before being eligible for the main experimental task. Experiments were conducted in each participant’s native language (English, French, Spanish, Korean, and Egyptian Arabic), instructions translated using GPT-4o. The repository accompanying this paper will include full task instructions and code to replicate all human experiments using the PsyNet framework [28].
@@ 3.2 Globally Representative Song Selection
To create a globally representative music dataset, we first used weekly YouTube top 100 music charts (year 2017-2023) from 59 countries, spanning six continents, ensuring representation from both Western and non-Western music. To ensure each country’s charts reflected its distinct popular music, we excluded any track appearing in more than one country’s chart. This left us with a country-exclusive pool of songs. From this pool, we sampled 20 songs per country, yielding 1,180 songs in total. Note that this diverse set is designed to capture a wide range of musical traditions and serve as a robust testbed for cross-cultural mood recognition. Each 15-second audio excerpt was trimmed from a random starting point in the full track, and then normalized at -5dB loudness.
In Stage 1 (Section 4.1), of the 1,180 full songs, we selected a subset of 180. For each country, we then added an additional 20 local songs (drawn from that country’s pool) to construct a country-specific subset of 200 songs (for participants from that country). This step ensured that local participants encounter enough music strongly tied to their background, allowing them to elicit culturally specific mood descriptors. In Stage 2 (Section 4.2), the entire set of 1,180 songs was rated by participants to capture a comprehensive cross-cultural perspective.
@@ 3.3 Model Evaluation
Leveraging our novel GlobalMood dataset, we evaluated several recent multimodal and multilingual models capable of music understanding. Specifically, we assessed Google’s Gemini models (1.5 Flash, 2.0 Flash, and the latest 2.5 Pro), a family of multimodal large language models capable of processing and reasoning across text and audio (but also image and video). We compared zero-shot and few-shot approaches, where the latter included 10 human-rated mood terms as examples.
Given that Gemini is closed-source, we also included CLAP (Contrastive Language-Audio Pretraining) [29] as an alternative, open-source model that learns joint audio-text embeddings. CLAP has demonstrated promise in MIR applications [30] and serve as the foundation for music-specific models like CLaMP [31]. Here, we conducted zero-shot evaluations through: (1) extracting audio embeddings from CLAP, (2) computing cosine similarities with text embeddings of mood terms, and (3) comparing these scores to human ratings.
We also fine-tuned CLAP on GlobalMood (train–test split = 1,000:180) to assess potential performance improvements. To preserve the continuous nature of our ratings, we represented each mood term in proportion to its mean rating (e.g., the term‘calm’, with a mean rating of 3.0, appeared three times in the text). This method retained the nuanced information in our soft labels rather than reducing them to binary categories. To improve generalizability, we created 10 augmented variations of each song through pitch shifting (range of $\pm$3 semitones), loudness adjustment (range of $\pm$15dB), and the addition of noise (amplitude of 0.005). Each augmented variant randomly included one or two of these modifications. The fine-tuned model checkpoints will also be publicly released.
Note that preliminary tests with recent multimodal models showed performance issues—Flamingo 2 [32] struggled with rating consistency and GPT-4o [33] failed to generate musical descriptions or ratings from audio alone—thus we excluded them from further analysis.
@@ 4.1.1 Tagging pipeline
Many existing studies on music emotions rely on pre-defined taxonomies or web-scraped data that offer limited linguistic diversity [8, 17]. To overcome this limitation, we employed a bottom-up, participant-driven tagging method [13, 14, 11]. Specifically, we asked participants in each country to complete independent ‘chains’ of iterative annotations on each of 200 representative sample songs (see Section 3.2 for details of the stimuli).
Figure 1A illustrates one such chain: (i) The first participant annotates the song using single-word mood or emotion tags in their native language; (ii) The second participant rates the relevance of these tags (1–5 scale), flag irrelevant tags (e.g., genre- or lyrics-related rather than mood), add new tags as necessary; (iii) The third participant sees all tags from earlier participants and repeats these steps; (iv) This iterative process continues through ten participants per chain, systematically refining and validating mood terms.
To mitigate potential bias from earlier participants’ contributions (i.e., priming effect), we conducted two parallel sets of annotation chains for each country.
4.1.2 Top emerging mood terms
Following the removal of tags flagged by more than two participants in a chain, our STEP-Tag process yielded an extensive, culturally specific lexicon of mood terms across languages: 644 unique terms in English, 528 in French, 870 in Spanish, 629 in Korean, and 283 in Arabic. To identify the most salient terms in each language, we calculated a composite score for every term by multiplying its frequency of occurrences across chains by its mean relevance rating. Higher scores indicate terms that participants frequently mentioned and consistently rated as highly relevant.
We consolidated closely related morphological variants (e.g., ‘happy’ and ‘happiness’ in English; gendered forms such as ‘joyeux’ and ‘joyeuse’ in French) manually with native speakers. This process was straightforward and unambiguous, and indeed, repeating this process with LLM (GPT-4o) provided nearly identical results. Figure 1B presents the resulting 20 highest-ranking mood tags per language, displaying both the original word and English translations to facilitate cross-cultural comparisons.

### v2

@@ Abstract
Human annotations of mood in music are essential for music generation and recommender systems. However, existing datasets predominantly focus on Western songs with terms derived from English, which may limit generalizability across diverse linguistic and cultural backgrounds. We introduce ‘GlobalMood’, a novel cross-cultural benchmark dataset comprising 1,180 songs sampled from 59 countries, with large-scale annotations collected from 2,519 individuals across five culturally and linguistically distinct locations: U.S., France, Mexico, S. Korea, and Egypt. Rather than imposing predefined emotion and mood categories, we implement a bottom-up, participant-driven approach to organically elicit culturally specific music-related emotion terms. We then recruit another pool of human participants to collect 988,925 ratings for these culture-specific descriptors. Our analysis confirms the presence of a valence-arousal structure shared across cultures, yet also reveals significant divergences in how certain emotion terms (despite being dictionary equivalents) are perceived cross-culturally. State-of-the-art multimodal models benefit substantially from fine-tuning on our cross-culturally balanced dataset, particularly in non-English contexts. Broadly, our findings inform the ongoing debate on the universality versus cultural specificity of emotional descriptors, and our methodology can contribute to other multimodal and cross-lingual research.
@@ 1 Introduction
Figure 1: Elicitation and refinement of music emotion terms through iterative participant chains. (A) Schematic illustration of the collaborative tagging process within a participant chain. Participants contribute new emotion-related word tags for each song, rate the relevance of existing tags, and can also flag irrelevant content, creating a dynamic refinement system. (B) Twenty most reliable emotion tags in each language, ranked by their tag scores. Y-axis labels display tags in their original language (left) and English translations (right).
@@
Music evokes diverse emotional responses in listeners, spanning a wide spectrum beyond basic emotional categories [1, 2]. A central challenge in Music Information Retrieval (MIR) is designing algorithms that can replicate this emotional sensitivity. This is crucial for building recommendation systems that align with listeners’ mood and context [3, 4, 5], and for generating music that resonates with individual preferences [6]. More broadly, understanding how music conveys emotion is a core question in the science of music [7, 8, 9]. To date, however, most algorithms have been trained on datasets derived from Western listeners and Western music, using taxonomies primarily based on English language (e.g., MIREX [10]).
@@
A significant challenge is creating cross-cultural models capable of handling non-Western music and emotion vocabularies beyond English. Addressing this challenge is essential to developing algorithms that accurately reflect global users’ preferences, including those whose musical tastes extend beyond the limited range of styles currently represented in training datasets. Moreover, without capturing culturally specific nuances of emotion, especially those difficult to translate, key aspects of musical meaning may be missed entirely. Direct dictionary translations of English terms may be insufficient, as terms describing emotions are deeply cultural and may lack exact equivalents [11, 12, 13, 14].
@@
To address these issues, we introduce ‘GlobalMood’, [^1] All code and data: https://github.com/harin-git/GlobalMood a new benchmark dataset designed to support culturally inclusive and linguistically diverse emotion and mood recognition in music. Our contribution innovates along three key dimensions: (i) the diversity of musical stimuli, drawn from 59 countries; (ii) the diversity of annotators, spanning five distinct regions (with plans to extens to over 20 languages and locations in future); (iii) a data-driven approach for collecting descriptors, generated organically by participants in their own language during the annotation process.
@@
Data were collected through two stages involving a total of 2,519 participants and 1,180 songs balanced evenly across 59 countries: In the first stage (Section 4.1; Figure 1), using a smaller subset of 200 songs, we employed our recently developed iterative task that combines open-ended elicitation with collective refinement [15, 16, 13]. Rather than asking listeners to choose from a fixed list of pre-defined emotion terms, we asked them to describe the perceived emotion conveyed in the music using free-text tags in their native language, and at the same time, rate the tags provided by previous listeners. This approach was key to uncovering emotion terms that would otherwise be overlooked by predefined, English-based taxonomies (such as ‘appeal/plead’ that appears in Korean only).
@@
In the second stage (Section 4.2; Figure 2), we selected the top 20 elicited terms per language and crowdsourced ratings for each tag across the entire set of 1,180 songs. This resulted in a total of 988,925 ratings, creating the most comprehensive open-source cross-cultural emotion annotation dataset in Music Emotion Recognition (MER) to date.
@@
We leveraged GlobalMood to test several recent multimodal and multilingual models (Gemini, CLAP) by evaluating their performance under zero-shot, few-shot, and fine-tuned scenarios (Section 4.3; Figure 3). Models trained only on English data performed poorly in some cultural contexts, but fine-tuning with our cross-cultural data greatly improved their performance in non-English settings. This highlights the critical importance of cross-cultural data in both training MER models and establishing appropriate benchmarks for their evaluation.
@@
Figure 2: Association between emotion terms across languages. (A) MDS visualization of the emotion terms based on mean ratings across the full song set. Terms positioned closer together exhibit similar rating patterns across songs, suggesting similar interpretations across languages. (B) Comparison of terms with direct translation equivalents across languages. The area size indicates the degree of semantic divergence despite apparent translation equivalence.
@@ 2 Related Works
2.1 Music Emotion and Mood Annotation Datasets
@@
Several datasets have been developed for MER systems with varying annotation approaches. [^2] Note that databases often extend the concept of emotion to include related constructs such as mood or feeling. Here we adopt this broader perspective, while acknowledging that subtle distinctions between them do exist. Early examples include the widely used MIREX 2007 mood dataset [10] with 240-250 Western songs in five mood clusters derived from AllMusic’s English tags (e.g., ‘passionate–rousing’, ‘wistful–bittersweet’), and CAL500 [17] with 500 Western pop/rock songs annotated using 18 English mood terms by U.S. undergraduate listeners. Over time, larger datasets appeared: the DEAM corpus (MediaEval ‘Emotion in Music’ dataset [18]) containing 2,058 song excerpts with continuous valence/arousal annotations; mood tags mined from large corpora of Spotify music playlists [19]; and the MTG-Jamendo dataset [20], which provides mood/theme tags for 18,486 songs. Notably, Jamendo’s tags were freely crowdsourced (56 unique mood labels), which introduced more label variety but still almost entirely in English.
@@
A common limitation across these datasets is their reliance on predefined English descriptors, many of which stem from Western music psychology (for an exception, see Strauss et al. [21]). For instance, the Geneva Emotional Music Scale (GEMS) defines 45 emotion descriptors (e.g., ‘joyful activation’) based on studies with European listeners [1], and this taxonomy has been used to annotate datasets like Emotify [22]. Similarly, the mood categories in MIREX and CAL500 were fixed in advance (drawn from AllMusic or prior literature) and presented to annotators as a closed set of options. Consequently, these top-down approaches restrict annotators to the moods the researchers envisioned, leaving any unlisted mood nuances uncaptured and undocumented.
@@
Acknowledging these limitations, recent research has begun exploring MER beyond the Western-centric scope. Hu et al. [23] examined mood annotations of K-pop songs provided by both Korean and American listeners. Their approach involved translating the original MIREX mood categories into Korean for local annotators. Although this allowed direct comparisons of mood classification between Korean and American listeners, it inherently restricted Korean annotations to terms originally defined within Western contexts.
@@
More recently, we compiled a balanced set of American, Brazilian, and Korean songs and gathered mood annotations across nine categories, where annotators rated songs both from their own and the other two countries [12]. We showed that certain mood terms like ‘energetic’ and ‘sad’ are highly consistent across cultures, while more abstract concepts like ‘love’ and ‘dreamy’ diverge considerably. Similar findings have been reported by other studies [14, 13], highlighting that when mood descriptors are imposed from one language onto another, important meanings can simply be ‘lost in translation’.
@@
In summary, while existing MER datasets and research have laid a solid groundwork, they remain limited by insufficient linguistic and cultural diversity. Because many are predominantly English-based and rely on top-down annotation strategies, they may overlook how people in other cultural contexts perceive emotion and mood in music.
@@ 2.2 Audio LLMs: the New Frontier in Music Tagging
Recent advances in multimodal large language models (LLMs) have opened promising avenues for downstream MIR tasks, including emotion recognition. These models combine the reasoning capabilities of LLMs with audio perception systems (audio LLMs), enabling more flexible and nuanced music understanding than traditional classification approaches [24, 25, 26].
Models like MuLan [27] and MERT [24] have demonstrated potential for zero-shot music emotion and mood classification by embedding audio and natural language descriptions in a shared semantic space. However, comprehensive benchmarks such as the MuChoMusic [25] highlight a crucial limitation: these models rely heavily on language modality and do not attend sufficiently to audio, often failing with more nuanced audio examples for downstream MIR tasks. This limitation could be particularly critical for non-Western music and non-English emotion descriptors, given that their training data are largely from Western contexts.
@@ 3.1 Participants
We recruited two independent sets of participants across the two stages of our data collection: Stage 1 for emotion term elicitation (N = 778; see Section 4.1) and Stage 2 for subsequent ratings on top 20 terms (N = 1,741; see Section 4.2). Participants had to be at least 18 years old, reside in the target country, and speak the target language as their primary language. Participants from the US were recruited through Prolific, while participants from the other four countries (France, Mexico, S. Korea, and Egypt) were recruited through the CINT platform. All participants provided informed consent under an approved protocol (see Section 7). Participants were instructed to wear headphones and had to pass a headphone screening task [29], and a language proficiency test [30] before being eligible for the main experimental task. Experiments were conducted in each participant’s native language (English, French, Spanish, Korean, and Egyptian Arabic), with instructions translated using GPT-4o. Code to replicate the experiment through the PsyNet framework [31] and all data are available at https://github.com/harin-git/GlobalMood
@@ 3.2 Globally Representative Song Selection
To create a globally representative music dataset, we used weekly YouTube top 100 music charts (year 2017-2023) from 59 countries, spanning six continents. To ensure each country’s charts reflected its distinct popular music, we excluded any track appearing in more than one country’s chart. This left us with a country-exclusive pool of songs. From this pool, we sampled 20 songs per country, yielding 1,180 songs in total. This diverse set is designed to capture a wide range of musical traditions and serve as a robust testbed for cross-cultural emotion recognition. Each 15-second audio excerpt was trimmed from a random starting point in the full track, and normalized at -5dB loudness.
@@ 3.3 Model Evaluation
We used the resulting GlobalMood dataset to evaluate several recent multimodal and multilingual models capable of music understanding. Specifically, we assessed Google’s Gemini models (1.5 Flash, 2.0 Flash, and the latest 2.5 Pro), a family of multimodal large language models capable of processing and reasoning across text and audio (but also image and video). [^3] Preliminary tests with other recent multimodal models showed performance issues—Flamingo 2 [32] struggled with rating consistency and GPT-4o [33] failed to generate musical descriptions or ratings from audio alone—thus we excluded them from further analysis. We compared zero-shot and few-shot approaches, where the latter included 10 human-rated emotion terms as examples.
Given that Gemini is closed-source, we also included CLAP (Contrastive Language-Audio Pretraining) [34] as an alternative, open-source model that learns joint audio-text embeddings. CLAP has demonstrated promise in MIR applications [35] and serves as the foundation for music-specific models like CLaMP [36]. Here, we conducted zero-shot evaluations through: (1) extracting audio embeddings from CLAP, (2) computing cosine similarities with text embeddings of emotion terms, and (3) comparing these scores to human ratings.
We also fine-tuned CLAP on GlobalMood (train–test split = 1,000:180) to assess potential performance improvements. To preserve the continuous nature of our ratings, we represented each term in proportion to its mean rating (e.g., the term ‘calm’, with a mean rating of 3.0, appeared three times in the text). This method retained the nuanced information in our soft labels rather than reducing them to binary categories. To improve generalizability, we created 10 augmented variations of each song through pitch shifting (range of $\pm$3 semitones), loudness adjustment (range of $\pm$15dB), and the addition of Gaussian noise (amplitude of 0.005). Each augmented variant randomly included one or two of these modifications.
@@ 4.1.1 Tagging pipeline
Many existing studies on music emotions rely on pre-defined taxonomies or web-scraped data that offer limited linguistic diversity [10, 19]. To overcome this limitation, we employed a bottom-up, participant-driven tagging method [15, 16, 13]. Specifically, we asked participants in each country to complete independent ‘chains’ of iterative annotations. A subsample of 200 songs from the 1,180 entire set was used as stimuli. This subsample consisted of 180 balanced songs across countries, with an additional 20 local songs drawn from the participating country’s pool. This was to ensure that local participants encounter enough music strongly tied to their background, allowing them to elicit culturally specific emotion descriptors.
Figure 1A illustrates one such chain: (i) The first participant annotates the song using single-word emotion tags in their native language; (ii) The second participant (from the same country) rates the relevance of these tags (1–5 scale), flag irrelevant tags (e.g., genre- or lyrics-related rather than emotion), add new tags as necessary; (iii) The third participant sees all tags from earlier participants and repeats these steps; (iv) This iterative process continues through ten participants per chain, systematically refining and validating emotion terms. In each country, we ran the entire elicitation experiment twice and aggregated the results to increase the diversity of responses from a larger pool of participants.
4.1.2 Top emerging terms
Following the removal of tags flagged by more than two participants in a chain, our STEP-Tag process yielded an extensive, culturally specific lexicon of emotion terms across languages (N unique terms: English = 644; French = 528; Spanish = 870; Korean = 629; Arabic = 283). To identify the most salient terms in each language, we calculated a composite score for every term by multiplying its frequency of occurrences across chains by its mean relevance rating. Higher scores indicate terms frequently mentioned and consistently rated as highly relevant.
We consolidated closely related morphological variants (e.g., ‘happy’ and ‘happiness’ in English; gendered forms such as ‘joyeux’ and ‘joyeuse’ in French) manually with native speakers. Figure 1B presents the resulting 20 highest-ranking tags per language, displaying both the original word and English translations to facilitate cross-cultural comparisons.
Despite being explicitly asked to provide emotion terms, participants often gave broader affective descriptors like moods or feelings (e.g., ‘soft’ and ‘festive’). This aligns with prior MIR literature that often includes both emotion and mood, and given its relevance for practical use, we did not enforce a strict distinction.